Find the value of "X" in a way that A= 12î + xj + 6k and B= 8î -3j -6k are perpendicular.

23. Encontrar el valor de x de forma que λ 12i + x) + 6E y B-81-3)-6E sean perpendiculares
Answer #1

Since,\underset{x}{\rightarrow}.\underset{y}{\rightarrow} = |\underset{x}{\rightarrow}|*|\underset{y}{\rightarrow}|*Cos(\Theta)

here \Theta = angle between vectors

if vectors are perpendicular then \Theta = 90 deg

So, Cos(\Theta) = 0

then, \underset{x}{\rightarrow} .\underset{y}{\rightarrow} = 0

therefore two vectors are perpendicular if there dot product is equals to zero,

here, \underset{A}{\rightarrow} = 12 i + x j + 6 k

\underset{B}{\rightarrow} = 8 i - 3 j - 6 k

Now, \underset{A}{\rightarrow} .\underset{B}{\rightarrow} = (12 i + x j + 6 k).(8 i - 3 j - 6 k)

also i.i = j.j = k.k = 1

i.j = j.k = k.i = 0

So,\underset{A}{\rightarrow}.\underset{B}{\rightarrow} = 12*8 + x*(-3) + 6*(-6)

\underset{A}{\rightarrow}.\underset{B}{\rightarrow} = 96 - 3x - 36

\underset{A}{\rightarrow}.\underset{B}{\rightarrow} = 60 - 3x

for perpendicular \underset{A}{\rightarrow} .\underset{B}{\rightarrow} = 0

60 - 3x = 0

3x = 60

x = 20

Please upvote.